How many hours would it take to go 12.8 miles if you were traveling at a rate of 163 cm/sec?
1
Expert's answer
2020-11-10T14:03:43-0500

"12.8mi=12.8\times1.6093=20.6km=2.06\times10^6cm"


"t=\frac{2.06\times10^6cm}{163cm/sec}\times\frac{1hour}{3600sec}=3.51hours"


Answer: 3.51 hours